40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Apr 24 16:01:55 2019, anchal, DailyProgress, FSS, Time series and spectrum analysis of RFout of SCavReflRFPD near resonance South_Cavity_Reflection_RFout_Time_Series_Analysis.pdfSouth_Cavity_Reflection_RFout_Spectrum_Analysis.pdf
    Reply  Fri Apr 26 14:34:03 2019, anchal, DailyProgress, FSS, Time series and spectrum analysis of RFout of SCavReflRFPD near resonance 
       Reply  Sun Apr 28 22:31:39 2019, rana, DailyProgress, FSS, RF beats 
          Reply  Tue Apr 30 18:07:49 2019, anchal, DailyProgress, FSS, Spectrum of RFout from South Cavity Reflection RFPD SCavReflRFoutSpecAnalysis.pdf
             Reply  Wed May 8 20:03:43 2019, anchal, DailyProgress, FSS, FSS PDH Error Signal on South Path is in good health SCavPDHError20190508.pdfCavity_Refl_RFPD_TI_After_Modifications.pdf
Message ID: 2338     Entry time: Wed May 8 20:03:43 2019     In reply to: 2331
Author: anchal 
Type: DailyProgress 
Category: FSS 
Subject: FSS PDH Error Signal on South Path is in good health 

Either by the mighty presence of Craig who came down to lab for few minutes or because of my recent updates in path alignment (see PSL:2335, PSL:2334 ), the PDH error signal in the south path looks almost as healthy as it ever was.

Attached is the data taken from oscilloscope with while scanning laser pzt at 3 Hz with 2 V peak-to-peak sinewave.

I measured modulation depth to be 0.373 by scanning the slow control voltage and reading powers in carrier and sidebands. Then I used the ratios of Bessel functions to estimate the modulation depth, which seems higher than expected.

The differences in the expected and the measured values good be due to many reasons like wrong modulation depth measurement, wrong responsivity (I used 0.75), wrong mixer loss and voltage divide estimation based on MAX333A datasheet values etc. So overall, this looks good enough that I can just move on. I am not sure how this actually happened.

Code and Data


Edit Mon May 13 16:40:21 2019 :

Corrected the transimpedance value of SN010 used. I was referring to a wrong value before. The have made the calculation by backtracing from the data in taken in CTN:2247, dividing by MAX4107 old gain (560/(0.5*113) +1), multiplying by the new gain (680/75 +1), dividing by old voltage division of (50/70) and multiplying by new voltage division of (50/100). The details are in this notebook.

Log of changes made to SN010, Schematic D980454-00;

1) Changed R1 to 680 Ohm and R2 to 75 Ohm. Replaced Max4107 with a new one.

2) Changed R6 to 50 Ohm. This changes the voltage division at output.

Log of changes made to SN009, Schematic D980454-00;

1) Changed R6 to 50 Ohm. This changes the voltage division at output.


Edit Wed May 15 19:47:42 2019:

See https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:rfpd for latest changes in RFPD.

Attachment 1: SCavPDHError20190508.pdf  200 kB  Uploaded Mon May 13 18:54:23 2019  | Hide | Hide all
SCavPDHError20190508.pdf
Attachment 2: Cavity_Refl_RFPD_TI_After_Modifications.pdf  1.646 MB  Uploaded Mon May 13 19:03:30 2019  | Hide | Hide all
Cavity_Refl_RFPD_TI_After_Modifications.pdf Cavity_Refl_RFPD_TI_After_Modifications.pdf Cavity_Refl_RFPD_TI_After_Modifications.pdf
ELOG V3.1.3-